40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 348     Entry time: Fri Apr 19 09:35:28 2019
Author: Joe 
Type: General 
Category:  
Subject: Adjusting cavity axis, re-alignment of OMC and locking 

[koji,philip, joe, liyuan, steven]

*still need to add photos to post*

PZT 11 was removed and inspected for so dust/dirt on the bottom of the prism. We saw a spot. We tried to remove this with acetone, but it stayed there. (Attachment 2, see the little white spec near the edge of the bottom surface of the prism)

current micrometer positions:

  • CM1: one closest to centre 11, close to edge 35 marking
  • CM2: both at 20 marking

Swapped PZT for PZT 22, cleaned the bottom and put it into position of CM1. We saw a low number of newton rings, so this is good.

We got a rough initial alignment by walking the beam with the periscope and PZT 22  mirrors. Once we saw a faint amount of transmission, we set up the wincam at the output. The reflected light from the cavity could also be seen to be flashing as the laser frequency was being modulated. 

Once it was roughly aligned, using the persicope we walked the beam until we got good 00 flashes. We checked the positions of the spots on the mirror with the beam card. This looked a lot better in the verticle direction (very near the centre) on both curved mirrors. We locked the cavity and contiued to align it better. This was done with the periscope until the DC error signal was about 0.6V. We switched to the fibre coupler after this. 

Once we were satisfied that he cavity was near where it would be really well aligned, we took some images of the spot positions. Using these we can work out which way to move the curved mirrors. Koji worked this out and drew some diagrams, we should attach them to this post. [Diagram: See Attachment 1 of ELOG OMC 350]

We made the corrections to the cavity mirrors

  • CM1: one closest to centre 11, close to edge 35+16 marking
  • CM2: I can't remember exactly, Koji created a diagram which would help explain this step [Diagram: See Attachment 2 of ELOG OMC 350]

The scatter from CM1 looked very small, it was hard to see with a viewer or CCD. We had to turn up the laser power by a factor of 3 to begin to see it, indicating that this is a good mirror.

Once this was done, the spot positions looked uch nearer the centre of each mirror. They look pitched 1mm too high, which might be because of the bottom surfaces of the prisms having a piece of dust on them? For now though it was good enough to try take the detuned locking FSR measurement and RFAM measurement. 

To see the higher order mode spacing, we misaligned them incoming beam in pitch and yaw so that the TM10 and TM01 modes were excited. The cavity transmission beam was aligned onto the photodiode such that we could make a transfer function measurement (i.e. shift the beam along the photodiode so that only half of the beam was on it, this maximises the amount of photocurrent).

attachment 1 shows the fitting of the detuned locking method for measuring FSR and cavity length/

I saved this data on my laptop. When I next edit this post (hopefully I will before monday, although I might be too tired from being a tourist in california...) I want to upload plots of the higher order mode spacing.

 

 

Attachment 1: FSR_Scan_Fitfsrdata.pdf  36 kB  Uploaded Mon Apr 22 10:07:13 2019  | Hide | Hide all
FSR_Scan_Fitfsrdata.pdf
Attachment 2: IMG_7679_cropped.jpg  87 kB  Uploaded Mon Apr 22 10:13:54 2019  | Hide | Hide all
IMG_7679_cropped.jpg
ELOG V3.1.3-